OpenStudy (anonymous):

Polar coordinates of a point are given. Find the rectangular coordinates of the point. (-5, -180°)

OpenStudy (dumbcow):

\[x = r \cos \theta = -5 \cos(-180) = 5\] \[y = r \sin \theta = -5 \sin (-180) = 0\]